You need $25,256 at the end of 8 years, and your only investment outlet is an 8 percent long-term certificate of deposit (compounded annually). With the certificate of deposit, you make an initial investment at the beginning of the first year. Use Appendix B and Appendix C for an approximate answer, but calculate your final answer using the formula and financial calculator methods.

  
a. What single payment could be made at the beginning of the first year to achieve this objective? (Do not round intermediate calculations. Round your final answer to 2 decimal places.)
  

Single payment made 


 


b. What amount could you pay at the end of each year annually for 8 years to achieve this same objective? (Do not round intermediate calculations. Round your final answer to 2 decimal places.)
  

amount to be made 


Please explain how to explain on financial calculator 
Answer #1

a)

PV = FV / (1 + i)^n

= 25,256 / (1.08)^8

=13,645.03

single payment = 13,645.03

b)

A = FV/ {[(1 + i)^n-1] / i}

= 25,256 / {[(1.08)^8 - 1 / .08}

= 2,374.44 (A= Annual payment, FV=Future value)

Amount to be paid =2,374.44